Application problem.
Find two negative consecutive even integers whose product is 120.

Found 2 solutions by Merciful, Alan3354:
Answer by Merciful(5)

n(n-2)=120
n^2 - 2n - 120 = 0
(n - 12)(n 10) = 0
n = -10 or n = 12 but 12 is N/A
The next negative even integer is -12
So the integers are -10 and -12.

Answer by Alan3354(69443)

Find two negative consecutive even integers whose product is 120.
---------
Try pairs of factors of 120:
1*120 NG
2*60 NG
etc